OpenStudy (anonymous):

what is the value of the expression? tan 2pie/3

OpenStudy (anonymous):

it is \[\frac{\sin(\frac{2\pi}{3})}{\cos(\frac{2\pi}{3})}\] so you need to know those two numbers to start

OpenStudy (anonymous):

and what do you do next

OpenStudy (anonymous):

find \(\sin(\frac{2\pi}{3})\) and \(\cos(\frac{2\pi}{3})\)

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet find the point on the unit circle corresponding to the angle \(\frac{2\pi}{3}\) the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

so wats the answer

OpenStudy (anonymous):

sqrt3?

OpenStudy (anonymous):

1 over sqrt3?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

none of those

OpenStudy (anonymous):

if you look at the cheat sheet you should see that at \(\frac{2\pi}{3}\) the first coordinate is \(-\frac{1}{2}\) and the second coordinate is \(\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

take the second coordinate, divide it by the first the 2's will cancel, but you have a minus sign there, so the correct answer is \(-\sqrt{3}\)